OpenStudy (timaashorty):

Can someone show me how to do this quickly? Use the law of sines to find the missing angle of the triangle. Find m ∠C to the nearest tenth/

OpenStudy (timaashorty):



OpenStudy (timaashorty):

A. 11.1 degrees Or B. 58.8 degrees.

OpenStudy (timaashorty):

@whpalmer4 @kropot72

OpenStudy (timaashorty):

@johnweldon1993

OpenStudy (johnweldon1993):

Ahh the law of sines lol \[\large \frac{sinA}{a} = \frac{sinB}{b} = \frac{sinC}{c}\]

OpenStudy (johnweldon1993):

So this means if we have an angle of A...and a side of a......along with at least 1 other angle...or 1 other side...we can solve for the missing (angle or side) that we need... So here We have angle A...and the corresponding side A We also have side c...so we can solve for angle C

OpenStudy (johnweldon1993):

\[\large \frac{sin(24^\circ)}{77} = \frac{sin(C)}{162}\]

OpenStudy (johnweldon1993):

Now we just need to isolate \(\large sin(C)\) which would look like...?

OpenStudy (timaashorty):

Sin24/77 x sin c/162

OpenStudy (timaashorty):

I don't understand the isolating part.

OpenStudy (timaashorty):

Do I just solve for C ?

OpenStudy (johnweldon1993):

That's alright :) so yes we have \[\large \frac{sin(24^\circ)}{77} = \frac{sin(C)}{162}\] If we cross multiply here (can only do this with ratios... this fraction = that fraction) we would have \[\large 162sin(24^\circ) = 77sin(C)\] does that make sense? Just multiply